A movie theater charges $5 for adults ticket and $2 for childs ticket. One Saturday the theater sold 785 tickets for a total of $3280.00. How many children's tickets were sold for the movie that Saturday?
I need the equations and steps for this problem. Thank you so much!!!

Answer by mananth(16949)

A movie theater charges $5 for adults ticket and $2 for childs ticket. One Saturday the theater sold 785 tickets for a total of $3280.00. How many children's tickets were sold for the movie that Saturday?
..
number of adult tickets sold be x
number of child tickets sold be y
..
x+y =785.....................1
5x+2y=3280...................2
...
multiply (1) by -5
-5x-5y=-3925
...
add this equation to(2)
5x-5x-5y+2y=-3925+3280
-3y=-645
/3
y = 215
the number of child tickets sold= 215
...
m.ananth@hotmail.ca